# On a B-field Transform of Generalized Complex Structures over Complex Tori

### v1

@@ 1.2 Motivation
First, let us recall the generalized complex structure induced from the complex structure on $X^{n}$ and B-field transforms of it. We assume that the period matrix of $X^{n}$ is given by a regular [^2] This regularity assumption is not necessary when we generally define an $n$-dimensional complex torus $X^{n}=\mathbb{C}^{n}/\mathbb{Z}^{n}\oplus T\mathbb{Z}^{n}$. This is necessary in order to define a mirror partner of $X^{n}$ (see [21]). matrix $T\in M(n;\mathbb{C})$ such that the imaginary part of $T$ is positive definite, i.e., $X^{n}=\mathbb{C}^{n}/\mathbb{Z}^{n}\oplus T\mathbb{Z}^{n}$, where $M(n;\mathbb{C})$ denotes the set of complex matrices of order $n$ (we use the notations $M(n;\mathbb{R})$ and $M(n;\mathbb{Z})$ in this sense). We sometimes decompose the matrix $T$ into its real part $X\in M(n;\mathbb{R})$ and its imaginary part $Y\in M(n;\mathbb{R})$: $T=X+\mathbf{i}Y$, where $\mathbf{i}:=\sqrt{-1}$. For this $X^{n}$, we can actually take an open covering $\{O_{i}\}_{i\in I}$ such that the open sets $O_{i}$ and their intersections are all contractible, and we locally express the complex coordinates of $O_{i}$ ($i\in I$) as $x+Ty$, where
@@
called the B-field, and define a alternating matrix $B\in M(2n;\mathbb{R})$ by
$B:=\left(\begin{array}[]{cc}B_{1}&B_{2}\\ -B_{2}^{t}&B_{3}\end{array}\right).$
@@
$\left(\begin{array}[]{cc}I_{n}&O\\ -B&I_{n}\end{array}\right)\left(\begin{array}[]{cc}I_{T}&O\\ O&-I_{T}^{t}\end{array}\right)\left(\begin{array}[]{cc}I_{n}&O\\ B&I_{n}\end{array}\right),$
where $I_{n}$ denotes the identity matrix of order $n$.
Now, we roughly explain the correspondence between gerby deformations of $X^{n}$ and B-field transforms of $\mathcal{I}_{T}$. In our setting, although we can construct a gerbe in the sense of Hitchin-Chatterjee as a family of smooth complex line bundles on $O_{i}\cap O_{j}$ ($i$, $j\in I$) (cf. [7, 13]), such smooth complex line bundles are all trivial since $O_{i}\cap O_{j}$ is contractible for each $i$, $j\in I$, namely, such a gerbe is trivial in fact. However, we can consider a “flat” connection in the sense of Hitchin-Chatterjee, i.e., 0-connection, and 1-connection compatible with it over this trivial gerbe. In particular, 1-connection is a 2-form (not a 1-form), and it can be regarded as the above B-field (1).
By concerning the above, in our previous work [24], we discuss the case of “$B_{2}$ type”, i.e., the case such that $B_{2}=\tau\in M(n;\mathbb{Z})$ [^3] Precisely speaking, we consider the case $B_{2}=\tau^{t}$ in [24]. and $B_{1}=B_{3}=O$. Actually, although we can consider the B-field transform of $\mathcal{I}_{T}$ in the case of “$B_{2}$ type” itself for $B_{2}=\tau\in M(n;\mathbb{R})$, the case $B_{2}=\tau\in M(n;\mathbb{Z})$ is particularly important in the sense that the corresponding deformation is closely related to a generalization of the $SL(2;\mathbb{Z})$-action (up to shifts) over $D^{b}(Coh(X^{1}))$ in the case of elliptic curves (cf. [29, 20]) to the higher dimensional case (we omit the details of this fact here), so we focus on the case $B_{2}=\tau\in M(n;\mathbb{Z})$ only in [24]. In this paper, as a sequel of our previous work [24], we investigate the B-field transform of $\mathcal{I}_{T}$ in the case of “$B_{1}$ type”, i.e., the case such that $B_{1}=\tau\in M(n;\mathbb{R})$ and $B_{2}=B_{3}=O$, including the corresponding deformation in the symplectic geometry side, and discuss the homological mirror symmetry for the mirror pair of the deformed tori on the object level.
On the other hand, if we focus on the case “$B_{3}$ type”, i.e., the case such that $B_{3}=\tau\in M(n;\mathbb{R})$ and $B_{1}=B_{2}=O$, then we need to consider the “$\beta$-field” transform depending on $\tau\in M(n;\mathbb{R})$ of the corresponding generalized complex structure in the symplectic geometry side, and it is closely related to a deformation quantization of the (complexified) symplectic torus $\check{X}^{n}$. This will be discussed in [25].
@@ 1.3 Related works
From the viewpoint of generalized complex geometry which is introduced by Hitchin and Gualtieri [14, 11], when we regard a complex manifold as a generalized complex manifold, as explained in subsection 1.2, it is thought that a B-field transform of the generalized complex structure is interpreted as a gerby deformation of the given complex manifold. This interpretation is mentioned in [11]. Applications of generalized complex geometry to the mirror symmetry are particularly well studied by Oren Ben-Bassat in [3, 4]. There, more precisely, he discusses the homological mirror symmetry for generalized complex manifolds under “adapted requirement” (see [3, Definition 4.2] [^4] Note that the definition of the mirror transform used in [3, 4] differs from the definition employed in this paper.). For example, in our setting, the generalized complex structure $\mathcal{I}_{T}$ induced from the complex structure on $X^{n}$ whose period matrix $T$ is given by $\mathbf{i}Y$ ($Y\in M(n;\mathbb{R})$ is positive definite) satisfies “adapted requirement”. Moreover, the generalized complex structure induced from the canonical complex structure on a (general) complex manifold is also a typical example which satisfies “adapted requirement”. In particular, relatively general B-field transforms of the generalized complex structure induced from the canonical complex structure on a complex manifold (B-field transforms treated in this paper are special cases of them) and the mirror dual counterparts are briefly explained in subsection 5.2 in [4] (under “adapted requirement”).
As examples other related works, in [17, 18] (see also [6]), Kapustin and Orlov study the effects of B-fields on the bounded derived categories of coherent sheaves over abelian varieties (we do not assume that $X^{n}$ is an abelian variety in this paper) and the Fukaya categories over given symplectic manifolds, including the formulation of the statement of the “deformed version” of the homological mirror symmetry conjecture. In particular, although we do not treat superconformal vertex algebras in this paper, in [17, 18], they examine properties of the bounded derived categories of twisted sheaves over the deformed abelian varieties from the viewpoint of superconformal vertex algebras.
@@ 1.4 Main results and the plan of this paper
We take an arbitrary alternating matrix $\tau\in M(n;\mathbb{R})$, and fix it. In this paper, we first define the gerby deformed $n$-dimensional complex torus $X_{\mathcal{G}_{\tau}}^{n}$ by regarding the B-field transform of $B_{1}$ type: $B_{1}=\tau\in M(n;\mathbb{R})$ for $\mathcal{I}_{T}$ as the twisting by the flat gerbe $\mathcal{G}_{\tau}$ in the sense of Hitchin-Chatterjee’s work [7, 13] (see also [12]) as mentioned in subsection 1.1, subsection 1.2 and subsection 1.3, where $\tau\in M(n;\mathbb{R})$ determines the flat connection (identified with the B-field of $B_{1}$ type essentially), and also consider its mirror partner $\check{X}_{\mathcal{G}_{\tau}}^{n}$ via generalized complex geometry. Next, we give the deformation of each object $E(s,\mathcal{L}^{\nabla})\to X^{n}$ as the twisted holomorphic line bundle with the twisted integrable connection $E(s,\mathcal{L}^{\nabla}(\tau))\to X_{\mathcal{G}_{\tau}}^{n}$ in the complex geometry side. Moreover, we prove that these objects $E(s,\mathcal{L}^{\nabla}(\tau))$ forms a dg-category $DG_{X_{\mathcal{G}_{\tau}}^{n}}$ in Theorem 5.5. In particular, we can regard this dg-category $DG_{X_{\mathcal{G}_{\tau}}^{n}}$ as a natural deformation of the dg-category $DG_{X^{n}}$ consisting of objects $E(s,\mathcal{L})$. Note that these deformations in the complex geometry side are successfully discussed for any $\tau\in M(n;\mathbb{R})$. On the other hand, the symplectic geometry side is a little complicated. More explicitly, although we may discuss the deformation of objects $(s,\mathcal{L}^{\nabla})$ according to the definition [9, Definition 1.1] (see also [17, 18] and so on) in the case $\tau\in M(n;\mathbb{Z})$, it becomes necessary to reconsider how to define objects which should be treated on the deformed complexified symplectic torus under the assumption $\tau\not\in M(n;\mathbb{Z})$. Also, in the symplectic geometry side, note that each Lagrangian section $s$ is not deformed since the symplectic structure on $\check{X}^{n}$ is preserved associated to the deformation from $X^{n}$ to $X_{\mathcal{G}_{\tau}}^{n}$, namely, each Lagrangian section $s$ is not deformed whether the condition $\tau\in M(n;\mathbb{Z})$ is assumed or not. By concerning the above problem, before considering the case $\tau\not\in M(n;\mathbb{Z})$, we focus on the case $\tau\in M(n;\mathbb{Z})$, and give the deformation of each pair $(s,\mathcal{L}^{\nabla})$ as the pair $(s,\mathcal{L}^{\nabla}(\tau))$ of the Lagrangian section $s$ of $\check{X}_{\mathcal{G}_{\tau}}^{n}\to\mathbb{R}^{n}/\mathbb{Z}^{n}$ and the smooth complex line bundle with the unitary connection $\mathcal{L}^{\nabla}(\tau)$ along it in the setting of the SYZ trivial special Lagrangian torus fibration $\check{X}_{\mathcal{G}_{\tau}}^{n}\to\mathbb{R}^{n}/\mathbb{Z}^{n}$ on the base space $\mathbb{R}^{n}/\mathbb{Z}^{n}$. Then, we can actually check that these deformed objects $E(s,\mathcal{L}^{\nabla}(\tau))$ and $(s,\mathcal{L}^{\nabla}(\tau))$ are mirror dual to each other, and this result is given in Proposition 5.7. Finally, we focus on the case $\tau\not\in M(n;\mathbb{Z})$. As a consequence, in this case, we can give the deformation of each unitary local system $\mathcal{L}^{\nabla}$ as a “twisted” smooth complex line bundles with a “twisted” unitary connection $\mathcal{L}^{\nabla}(\tau)$, and actually, an analogue of Proposition 5.7 also holds under the assumption $\tau\not\in M(n;\mathbb{Z})$. This result is given in Proposition 6.2. These imply that our deformations discussed in this paper seem to be natural from the viewpoint of the homological mirror symmetry for $(X_{\mathcal{G}_{\tau}}^{n},\check{X}_{\mathcal{G}_{\tau}}^{n})$.
@@
This paper is organized as follows. In section 2, we recall the definition of gerbes in the sense of Hitchin-Chatterjee briefly. In section 3, we construct a mirror partner $\check{X}^{n}$ of a given $n$-dimensional complex torus $X^{n}$ via generalized complex geometry. In section 4, we define the deformation $X_{\mathcal{G}_{\tau}}^{n}$ of $X^{n}$ by $\mathcal{G}_{\tau}$ as the generalized complex torus whose generalized complex structure is given by the B-field transform of $B_{1}$ type: $B_{1}=\tau\in M(n;\mathbb{R})$ for $\mathcal{I}_{T}$, and construct its mirror partner $\check{X}_{\mathcal{G}_{\tau}}^{n}$. Main discussions are provided in section 5 and section 6. The purpose of section 5 is to discuss the deformations on both sides associated to the deformation from $(X^{n},\check{X}^{n})$ to $(X_{\mathcal{G}_{\tau}}^{n},\check{X}_{\mathcal{G}_{\tau}}^{n})$ by focusing mainly on the case $\tau\in M(n;\mathbb{Z})$. In subsection 5.1, we recall the correspondence between holomorphic line bundles with integrable connections $E(s,\mathcal{L}^{\nabla})$ and pairs $(s,\mathcal{L}^{\nabla})$ of Lagrangian sections $s$ of $\check{X}^{n}\to\mathbb{R}^{n}/\mathbb{Z}^{n}$ with unitary local systems $\mathcal{L}^{\nabla}$ along them. In subsection 5.2, we discuss deformations $E(s,\mathcal{L}^{\nabla}(\tau))$ of objects $E(s,\mathcal{L}^{\nabla})$, and prove that these objects $E(s,\mathcal{L}^{\nabla}(\tau))$ naturally forms a dg-category $DG_{X_{\mathcal{G}_{\tau}}^{n}}$ in Theorem 5.5. In subsection 5.3, we discuss deformations $(s,\mathcal{L}^{\nabla}(\tau))$ of objects $(s,\mathcal{L}^{\nabla})$. Moreover, for given mirror dual objects $E(s,\mathcal{L}^{\nabla})$ and $(s,\mathcal{L}^{\nabla})$, we confirm that the deformations of them are also mirror dual to each other. This result is given in Proposition 5.7. In section 6, we consider the case $\tau\not\in M(n;\mathbb{Z})$. In subsection 6.1, we explain the problem which should be treated, and give the outline of an idea to solve the problem. In subsection 6.2, we describe the details of the idea which is mentioned in subsection 6.1, namely, we discuss deformations of objects $(s,\mathcal{L}^{\nabla})$ compatible with the counterpart in the mirror dual complex geometry side under the assumption $\tau\not\in M(n;\mathbb{Z})$. In particular, an analogue of Proposition 5.7 is given in Proposition 6.2.

### v2

@@ 1.2 Motivation
First, let us recall the generalized complex structure induced from the complex structure on $X^{n}$ and B-field transforms of it. We assume that the period matrix of $X^{n}$ is given by a regular [^2] This regularity assumption is not necessary when we define an $n$-dimensional complex torus $X^{n}=\mathbb{C}^{n}/\mathbb{Z}^{n}\oplus T\mathbb{Z}^{n}$. This is necessary in order to define a mirror partner of $X^{n}$ (see [22]). matrix $T\in M(n;\mathbb{C})$ such that the imaginary part of $T$ is positive definite, i.e., $X^{n}=\mathbb{C}^{n}/\mathbb{Z}^{n}\oplus T\mathbb{Z}^{n}$, where $M(n;\mathbb{C})$ denotes the set of complex matrices of order $n$ (we use the notations $M(n;\mathbb{R})$ and $M(n;\mathbb{Z})$ in this sense). We sometimes decompose the matrix $T$ into its real part $X\in M(n;\mathbb{R})$ and its imaginary part $Y\in M(n;\mathbb{R})$: $T=X+\mathbf{i}Y$, where $\mathbf{i}:=\sqrt{-1}$. For this $X^{n}$, we can take an open covering $\{O_{i}\}_{i\in I}$ such that the open sets $O_{i}$ and their intersections are all contractible, and we locally express the complex coordinates of $O_{i}$ ($i\in I$) as $x+Ty$, where
@@
called a B-field, and define an alternating matrix $B\in M(2n;\mathbb{R})$ by
$B=\left(\begin{array}[]{cc}B_{1}&B_{2}\\ -B_{2}^{t}&B_{3}\end{array}\right).$
@@
$\left(\begin{array}[]{cc}I_{2n}&O\\ -B&I_{2n}\end{array}\right)\left(\begin{array}[]{cc}I_{T}&O\\ O&-I_{T}^{t}\end{array}\right)\left(\begin{array}[]{cc}I_{2n}&O\\ B&I_{2n}\end{array}\right),$
where $I_{2n}$ denotes the identity matrix of order $2n$.
Now, we roughly explain the correspondence between gerby deformations of $X^{n}$ and B-field transforms of $\mathcal{I}_{T}$. In our setting, although we can construct a gerbe in the sense of Hitchin-Chatterjee as a family of smooth complex line bundles on $O_{i}\cap O_{j}$ ($i$, $j\in I$) (cf. [8, 14]), such smooth complex line bundles are all trivial since $O_{i}\cap O_{j}$ is contractible for each $i$, $j\in I$. This fact indicates that such a gerbe is trivial. However, we can consider a “flat” connection in the sense of Hitchin-Chatterjee, i.e., 0-connection, and 1-connection which is compatible with it over this trivial gerbe. In particular, 1-connection is a 2-form (not a 1-form), and it can be regarded as the above B-field (1).
By concerning the above, in our previous work [25], we discuss the case of “$B_{2}$ type”, i.e., the case such that $B_{2}=\tau\in M(n;\mathbb{Z})$ [^3] Precisely speaking, we consider the case $B_{2}=\tau^{t}$ in [25]. and $B_{1}=B_{3}=O$. Actually, although we can consider the B-field transform of $\mathcal{I}_{T}$ in the case of “$B_{2}$ type” itself for $B_{2}=\tau\in M(n;\mathbb{R})$, the case $B_{2}=\tau\in M(n;\mathbb{Z})$ is particularly important in the sense that the corresponding deformation is closely related to a generalization of the $SL(2;\mathbb{Z})$-action (up to shifts) over $D^{b}(Coh(X^{1}))$ in the case of elliptic curves (cf. [30, 21]) to the higher dimensional case (we omit the details of this fact here). Therefore, we focus on the case $B_{2}=\tau\in M(n;\mathbb{Z})$ only in [25]. In this paper, as a sequel of our previous work [25], we investigate the B-field transform of $\mathcal{I}_{T}$ in the case of “$B_{1}$ type”, i.e., the case such that $B_{1}=\tau\in M(n;\mathbb{R})$ and $B_{2}=B_{3}=O$, including the corresponding deformation in the symplectic geometry side, and discuss the homological mirror symmetry for the mirror pair of the deformed tori on the object level.
On the other hand, if we focus on the case “$B_{3}$ type”, i.e., the case such that $B_{3}=\tau\in M(n;\mathbb{R})$ and $B_{1}=B_{2}=O$, then we need to consider the “$\beta$-field” transform depending on $\tau\in M(n;\mathbb{R})$ of the corresponding generalized complex structure in the symplectic geometry side. In general, it is closely related to a deformation quantization of the (complexified) symplectic torus $\check{X}^{n}$. This will be discussed in [26].
@@ 1.3 Related works
From the viewpoint of generalized complex geometry which is introduced by Hitchin and Gualtieri [15, 12], when we regard a complex manifold as a generalized complex manifold, as explained in subsection 1.2, it is thought that a B-field transform of the generalized complex structure is interpreted as a gerby deformation of the given complex manifold. This interpretation is mentioned in [12]. Applications of generalized complex geometry to the mirror symmetry are particularly well studied by Oren Ben-Bassat in [3, 4]. There, more precisely, he discusses the homological mirror symmetry for generalized complex manifolds under “adapted requirement” (see [3, Definition 4.2]) [^4] Note that the definition of the mirror transform used in [3, 4] differs from the definition employed in this paper.. For example, in our setting, the generalized complex structure $\mathcal{I}_{T}$ induced from the complex structure on $X^{n}$ whose period matrix $T$ is given by $\mathbf{i}Y$ ($Y\in M(n;\mathbb{R})$ is positive definite) satisfies “adapted requirement”. Moreover, the generalized complex structure induced from the canonical complex structure on a (general) complex manifold is also a typical example which satisfies “adapted requirement”. In particular, relatively general B-field transforms of the generalized complex structure induced from the canonical complex structure on a complex manifold (B-field transforms treated in this paper are special cases of them) and the mirror dual counterparts are briefly explained in subsection 5.2 in [4] (under “adapted requirement”).
As examples of other related works, in [18, 19] (see also [7]), Kapustin and Orlov study the effects of B-fields on the bounded derived categories of coherent sheaves over abelian varieties (we do not assume that $X^{n}$ is an abelian variety in this paper) and the Fukaya categories over given symplectic manifolds, including the formulation of the statement of the “deformed version” of the homological mirror symmetry conjecture. In particular, although we do not treat superconformal vertex algebras in this paper, in [18, 19], they examine properties of the bounded derived categories of twisted sheaves over the deformed abelian varieties from the viewpoint of superconformal vertex algebras.
@@ 1.4 Main results and the plan of this paper
We take an arbitrary alternating matrix $\tau\in M(n;\mathbb{R})$, and fix it. In this paper, first, we define the gerby deformed $n$-dimensional complex torus $X_{\mathcal{G}_{\tau}}^{n}$ by regarding the B-field transform of $B_{1}$ type: $B_{1}=\tau\in M(n;\mathbb{R})$ for $\mathcal{I}_{T}$ as the twisting by the flat gerbe $\mathcal{G}_{\tau}$ in the sense of Hitchin-Chatterjee’s work [8, 14] (see also [13]), where $\tau\in M(n;\mathbb{R})$ determines the flat connection (identified with the B-field of $B_{1}$ type essentially), and also consider its mirror partner $\check{X}_{\mathcal{G}_{\tau}}^{n}$ via generalized complex geometry. Next, we give the deformation of each object $E(s,\mathcal{L}^{\nabla})\to X^{n}$ as the twisted holomorphic line bundle with the twisted integrable connection $E(s,\mathcal{L}^{\nabla}(\tau))\to X_{\mathcal{G}_{\tau}}^{n}$ in the complex geometry side [^5] Although it seems natural to write $E(s,\mathcal{L}^{\nabla})(\tau)$ instead of $E(s,\mathcal{L}^{\nabla}(\tau))$, we employ the notation $E(s,\mathcal{L}^{\nabla}(\tau))$ by concerning deformations in the symplectic geometry side.. Moreover, we prove that these objects $E(s,\mathcal{L}^{\nabla}(\tau))$ forms a dg-category $DG_{X_{\mathcal{G}_{\tau}}^{n}}$ in Theorem 5.7. In particular, we can regard this dg-category $DG_{X_{\mathcal{G}_{\tau}}^{n}}$ as a natural deformation of the dg-category $DG_{X^{n}}$ consisting of objects $E(s,\mathcal{L})$. On the other hand, the symplectic geometry side is a little complicated. More explicitly, although we can discuss the deformation of objects $(s,\mathcal{L}^{\nabla})$ according to the definition [10, Definition 1.1] (see also [18, 19] and so on) in the case $\tau\in M(n;\mathbb{Z})$, it becomes necessary to reconsider how to define objects which should be treated on the deformed complexified symplectic torus under the assumption $\tau\not\in M(n;\mathbb{Z})$. Also, in the symplectic geometry side, note that each Lagrangian section $s$ is not deformed since the symplectic structure on $\check{X}^{n}$ is preserved associated to the deformation from $X^{n}$ to $X_{\mathcal{G}_{\tau}}^{n}$. By concerning the above problem, we decompose $\tau$ into its integer part $\tau_{\mathbb{Z}}\in M(n;\mathbb{Z})$ and its fractional part $\tau_{\rm frac}\in M(n;\mathbb{R})$:
$\tau=\tau_{\mathbb{Z}}+\tau_{\rm frac}.$
Under this setting, we propose giving the deformation of each unitary local system $\mathcal{L}^{\nabla}$ as a “twisted” smooth complex line bundles with a “twisted” connection $\mathcal{L}^{\nabla}(\tau)$. Here, the twisting of $\mathcal{L}^{\nabla}$ is defined by using $\tau_{\rm frac}$, and at least, $\tau_{\mathbb{Z}}$ determines the first Chern character of $\mathcal{L}^{\nabla}(\tau)$ in the case $\tau\in M(n;\mathbb{Z})$, i.e., $\tau_{\rm frac}=O$. Then, we can check that these deformed objects $E(s,\mathcal{L}^{\nabla}(\tau))$ and $(s,\mathcal{L}^{\nabla}(\tau))$ are mirror dual to each other, and this result is given in Proposition 5.10. These imply that our deformations discussed in this paper seem to be natural from the viewpoint of the homological mirror symmetry for $(X_{\mathcal{G}_{\tau}}^{n},\check{X}_{\mathcal{G}_{\tau}}^{n})$.
@@
This paper is organized as follows. In section 2, we recall the definition of gerbes in the sense of Hitchin-Chatterjee briefly. In section 3, we construct a mirror partner $\check{X}^{n}$ of a given $n$-dimensional complex torus $X^{n}$ via generalized complex geometry. In section 4, we define the deformation $X_{\mathcal{G}_{\tau}}^{n}$ of $X^{n}$ by $\mathcal{G}_{\tau}$ as the generalized complex torus whose generalized complex structure is given by the B-field transform of $B_{1}$ type: $B_{1}=\tau\in M(n;\mathbb{R})$ for $\mathcal{I}_{T}$, and construct its mirror partner $\check{X}_{\mathcal{G}_{\tau}}^{n}$. Main discussions are provided in section 5. The purpose of section 5 is to discuss the deformations on both sides associated to the deformation from $(X^{n},\check{X}^{n})$ to $(X_{\mathcal{G}_{\tau}}^{n},\check{X}_{\mathcal{G}_{\tau}}^{n})$. In subsection 5.1, we recall the correspondence between holomorphic line bundles with integrable connections $E(s,\mathcal{L}^{\nabla})$ and pairs $(s,\mathcal{L}^{\nabla})$, i.e., Lagrangian sections $s$ of $\check{X}^{n}\to\mathbb{R}^{n}/\mathbb{Z}^{n}$ with unitary local systems $\mathcal{L}^{\nabla}$ along them. In subsection 5.2, we discuss deformations $E(s,\mathcal{L}^{\nabla}(\tau))$ of objects $E(s,\mathcal{L}^{\nabla})$, and prove that these objects $E(s,\mathcal{L}^{\nabla}(\tau))$ naturally forms a dg-category $DG_{X_{\mathcal{G}_{\tau}}^{n}}$ in Theorem 5.7. In subsection 5.3, we explain the problem which should be treated, and give the outline of an idea to solve the problem. In subsection 5.4, we describe the details of the idea which is mentioned in subsection 5.3, namely, we discuss deformations of objects $(s,\mathcal{L}^{\nabla})$ which are compatible with the counterpart in the mirror dual complex geometry side by focusing the division $\tau=\tau_{\mathbb{Z}}+\tau_{\rm frac}$. As a consequence, for given mirror dual objects $E(s,\mathcal{L}^{\nabla})$ and $(s,\mathcal{L}^{\nabla})$, we can confirm that the deformations of them are also mirror dual to each other. This result is given in Proposition 5.10. On the other hand, in general, the choice of a division of $\tau$ is not unique. For instance, when we focus on the trivial division $\tau=O+\tau$, we can also consider the twisting of $\mathcal{L}^{\nabla}$ associated to $\tau$. However, it can be expected that our proposal explained in this paper does not depend on the choice of such a division of $\tau$. We also describe this perspective in subsection 5.4.